9, A random sample is obtained from a population with variance = 400 and the sample mean is computed to be 70, Consider the null hypothesis Ho: μ = 80 versus the alternative H1: Ho: μ < 80. Compute the p-value. If n 32, the p-value is
Answer #1

Q.9) Given that, sample size ( n ) = 32

sample mean (\bar x) = 70

population variance (\sigma^2) = 400

population standard deviation (σ) V400-20

The null and the alternative hypotheses are,

Ho : μ = 80 versus H1 : μ < 80

Test statistic is,

Z = \frac { 70-80}{\frac { 20}{\sqrt { 32}}}= -2.83

Since, it is left-tailed test,

p-value = P(Z < -2.83) = 0.0023

p-value = 0.0023